OpenStudy (anonymous):

factor completely 16t^2+49-56t

OpenStudy (anonymous):

16t^2+49-56t = 16t^2-56t+49 = (4t)^2 - 2(4t)(7) + 7^2 = (4t - 7)^2 Basically, we rearrange the terms and apply the identity (a-b)^2 = a^2 - 2ab +b^2 here. Hope it helps.